# On residual finiteness of graphs of free groups with cyclic edge groups

### v1

@@ Abstract.
We characterize which groups splitting as finite graphs of free groups with cyclic edge groups are residually finite. Such a group $G$ is residually finite if and only if it does not contain a very unbalanced element, an element satisfying a small set of group equations and non-equations. From a presentation of $G$, we construct a finite labeled graph $\Gamma$, and show that the existence of very unbalanced elements in $G$ is equivalent to an easily-detectable property of this graph. This provides a simple algorithm deciding whether $G$ is residually finite. This characterization confirms a conjecture of Wise from [Wis00].
@@ 1. Introduction
Previously, Wise characterized the subgroup separable groups in this family, a much stronger property than residual finiteness [Wis00]. He proved the fundamental group of a graph of free groups with cyclic edge groups is subgroup separable if and only if there exists no unbalanced element, an infinite-order $g$ such that $g^{n}$ and $g^{m}$ are conjugate for $|n|$, $|m|$ distinct. We prove the analogous characterization of residual finiteness.
Definition.
A very unbalanced element in a group $G$ is an infinite-order $g\in G$ satisfying the following for some $h\in G$, $m,n\in\mathbb{Z}$, with $1$, $|m|$, and $|n|$ distinct:
  • $hg^{m}h^{-1}=g^{n}$
  • $hg^{m/k}h^{-1}\neq g^{n/k}$ for any $k>1$ dividing both $m$ and $n$.
Example 1.1.
Let $p,q$ be non-zero integers. The Baumslag-Solitar group $BS(p,q)$ is defined by the presentation $\langle a,t\mid ta^{p}t^{-1}=a^{q}\rangle$. If $1$, $|p|$, and $|q|$ are distinct, then the element $a\in BS(p,q)$ is very unbalanced. This is a consequence of Britton’s Lemma (see Lemma 4.4). If $1$, $|p|$, and $|q|$ are not distinct, then $BS(p,q)$ is residually finite (see [Mol18]), and we will prove (Proposition 3.5) that no very unbalanced elements exist in that case.
@@
Suppose $G$ splits as a finite graph of free groups with cyclic edge groups. Then $G$ is residually finite if and only if $G$ contains no very unbalanced elements.
The existence of a very unbalanced element is easily detectable from the graph of groups splitting of $G$ (Remark 4.9). In particular, there exists an algorithm which either confirms $G$ is residually finite or outputs a very unbalanced element.
@@
In Section 3 we define and discuss very unbalanced group elements. We show that in general a residually finite group can contain very unbalanced elements (Remark 3.6), but a residually finite group splitting as a graph of free groups cannot contain them (Proposition 3.5). In other words, when discussing graphs of free groups, very unbalanced elements serve as a certificate of non-residual finiteness. This provides one direction of Theorem A.
@@ 2.4. Small-cancellation theory
3. Very unbalanced elements
Recall that [Wis00] identified the presence of unbalanced elements as an obstruction for a group $G$ to be subgroup separable (i.e. LERF). An unbalanced element is an element $g\in G$ such that there exists $h\in G$, $n,m\in\mathbb{Z}$ with $|n|\neq|m|$ and $hg^{m}h^{-1}=g^{n}$. As we consider the weaker property of residual finiteness, we need a stronger obstruction.
@@ 3. Very unbalanced elements
In a group $G$, an element $g\in G$ is very unbalanced if there exists $h\in G$, $m,n\in\mathbb{Z}$, with $1$, $|m|$, and $|n|$ distinct, such that the following hold:
  • $g$ has infinite order
@@
  • $hg^{m/k}h^{-1}\neq g^{n/k}$ for any $k>1$ dividing both $m$ and $n$.
It clearly follows from the definition that being very unbalanced is a conjugacy class invariant. Fix an arbitrary group $G$.
@@
Let $g\in G$ be very unbalanced for $h,m,n$ and let $r\in G$ be a root of $g$ in the sense that $r^{L}=g$ for some $L\in\mathbb{Z}$. Then there exists an integer $l\geq 1$ such that $r$ is very unbalanced for $h,lm,ln$.
@@
Let $L$ be such that $g=r^{L}$. Since $g$ has infinite order, $r$ has infinite order. Let $l\geq 1$ dividing $L$ be minimal satisfying $hr^{lm}h^{-1}=r^{ln}$. Clearly, $1,|lm|,|ln|$ are distinct.
Assume for some $k>1$ dividing $lm$ and $ln$ that $hr^{lm/k}h^{-1}=r^{ln/k}$ holds. Set $k^{\prime}=k/gcd(l,k)$ and take both sides of the previous equation to the power $gcd(l,k)L/l$ to get $hr^{Lm/k^{\prime}}h^{-1}=r^{Ln/k^{\prime}}$. Substituting, we have $hg^{m/k^{\prime}}h^{-1}=g^{n/k^{\prime}}$. Hence $k^{\prime}=1$, $k$ divides $l$, and $l/k$ contradicts the minimality of $l$. ∎
The proof of the following lemma appears in [Mol18], though the lemma is not stated directly.
@@
Let $\varphi$ be a homomorphism from $G$ to a finite group. Let $g,h\in G$, $m,n\in\mathbb{Z}$ and $d=gcd(m,n)$. If $hg^{m}h^{-1}=g^{n}$, then the element $\varphi([hg^{d}h^{-1},g])$ is trivial.
@@
Since $\varphi(hgh^{-1})$ and $\varphi(g)$ are conjugate, they have the same order $\omega$. By assumption, $\varphi(hgh^{-1})^{m}$ and $\varphi(g)^{n}$ are equal and thus have the same order $\omega/gcd(\omega,m)=\omega/gcd(\omega,n)$. Thus $gcd(\omega,m)=gcd(\omega,n)$, so $gcd(\omega,m)$ divides $d$. The elements $m$ and $gcd(\omega,m)$ generate the same cyclic subgroup of $\mathbb{Z}/\omega\mathbb{Z}$, so we have $\varphi(hgh^{-1})^{d}\in\left\langle\varphi(hgh^{-1})^{gcd(\omega,m)}\right\rangle=\left\langle\varphi(hgh^{-1})^{m}\right\rangle=\left\langle\varphi(g)^{n}\right\rangle\subseteq\left\langle\varphi(g)\right\rangle$, and the lemma follows.
The following lemma contains standard, well-known results about free group elements.
@@
Any non-trivial element $g$ in a free group $F$ has a maximal root $r$ so that $g$ is a power of $r$ and $r$ is not a proper power. This root is unique up to inversion.
Moreover, given two non-trivial elements $g,h$ in $F$, the following are equivalent:
  1. (1)
    $g$ and $h$ commute
  2. (2)
    $g$ and $h$ have a common power
  3. (3)
    $g$ and $h$ are powers of the same element
Proof.
Consider a Cayley tree for $F$. If $g\in F$ is non-trivial, the subgroup $S_{g}<F$ of elements acting by translations on the axis of $g$ in the tree is infinite cyclic, as $S_{g}$ is a subgroup of the cyclic group of orientation-preserving automorphisms of a bi-infinite line graph. By uniqueness of the axis, any element with a power equal to $g$ stabilizes the axis of $g$. This means that the only two maximal roots of $g$ are the generators of $S_{g}$.
$(3)\Rightarrow(1)$ and $(3)\Rightarrow(2)$ are clear. If $(1)$ or $(2)$ is satisfied and $g,h$ are non-trivial, some power of $h$ preserves the axis for $g$ in a Cayley tree. By uniqueness of the axis of $h$, both $g$ and $h$ are powers of a maximal root of $g$. ∎
Proposition 3.5.
Assume that $G$ splits as a graph of free groups. If $G$ is residually finite, $G$ contains no very unbalanced element.
Here we make no finiteness assumptions on the splitting or the rank of the free groups. We make no assumptions on the edge groups of the splitting either.
Proof.
[…]
As conjugates, the elements $g^{m}$ and $g^{n}$ have the same translation length in the Bass-Serre tree of the graph of groups, yet $|m|\neq|n|$, hence $g$ is elliptic, stabilizing a vertex $v$. Assume that $h$ acts elliptically as well. Then $h$ fixes a vertex $w$ on the (possibly trivial) geodesic from $v$ to $hv$. However, the element $g^{n}=hg^{m}h^{-1}$ fixes both $v$ and $hv$, hence it must fix $w$. Thus the free group $Stab(w)$ contains $h,g^{m},g^{n}$, and $g^{d}\in\left\langle g^{m},g^{n}\right\rangle$. In a Cayley tree for $Stab(w)$, the elements $g^{m}$ and $g^{n}$ are powers of $g^{d}$ with the same translation lengths. This is a contradiction as $g$ is of infinite order and $|n|<|m|$. Thus, moving forward, we suppose that $h$ acts loxodromically on the Bass-Serre tree. Let $\alpha_{h}$ be its axis, with ends denoted $h^{\pm\infty}$.
Assume first that $Fix(g)$ does not contain the end $h^{+\infty}$, and let $x$ be the vertex of $Fix(g)$ closest to $h^{+\infty}$. More precisely:
  • If $Fix(g)$ doesn’t intersect $\alpha_{h}$, let $x$ be the vertex of $Fix(g)$ closest to $\alpha_{h}$. See Figure 1(a).
  • Otherwise, $Fix(g)$ intersects $\alpha_{h}$ along an interval that does not have $h^{+\infty}$ as an endpoint, let $x$ be the endpoint of that interval closest to $h^{+\infty}$. See Figure 1(b).
Let $r$ be a maximal root of $g$ in the free group $Stab(x)$, so that $x\in Fix(r)$. Since $Fix(r)\subseteq Fix(g)$, $Fix(r)$ does not contain $h^{+\infty}$, and $x$ is the point of $Fix(r)$ closest to $h^{+\infty}$. Observe that, in particular, the geodesic segment $\gamma$ joining $x$ to $hx$ only intersects $Fix(r)$ at $x$.
Figure 1. Different cases for the position of the point $x$, depending on whether $Fix(g)$ intersects $\alpha_{h}$ or not.
By Lemma 3.2, $r$ is very unbalanced for $h$ and some $lm,ln$. By Lemma 3.3 and residual finiteness of $G$, $[hr^{ld}h^{-1},r]$ is trivial in $G$. Hence $Fix(r)$ and $Fix(hr^{ld}h^{-1})$ intersect at some vertex $y$. The set $Fix(hr^{ld}h^{-1})$ is convex and contains both $hx$ and $y$. Thus it contains the geodesic segment $\gamma^{\prime}$ joining these two points. Since $y\in Fix(r)$, $\gamma^{\prime}$ contains $\gamma$. This proves that $x$ belongs to $Fix(hr^{ld}h^{-1})$. Therefore, $r$ and $hr^{ld}h^{-1}$ both lie in the free group $Stab(x)$.
Since $r$ and $hr^{ld}h^{-1}$ are commuting elements of a free group and $r$ is not a proper power, there exists $p\in\mathbb{Z}$ such that $hr^{ld}h^{-1}=r^{p}$ by Lemma 3.4. Then $r^{ln}=(hr^{ld}h^{-1})^{m/d}=r^{pm/d}$ and $ln=pm/d$, since $r$ has infinite order. Setting $k=m/d=ln/p$, the equation $hr^{lm/k}h^{-1}=r^{ln/k}$ holds, and because $r$ is very unbalanced, $|k|=1$ and thus $d=|m|$, a contradiction with the assumption that $1<n<|m|$.
Assume now that $Fix(g)$ contains the end $h^{+\infty}$. Let $x$ be a vertex in $\alpha_{h}\cap Fix(g)$, and $r$ a maximal root of $g$ in $Stab(x)$. As before, $r$ is very unbalanced for $h$ and some $lm,ln$. If $Fix(r)$ does not contain the end $h^{+\infty}$ the previous argument applied to $r$ instead of $g$ yields a contradiction. Otherwise, $Fix(r)$ contains $hx$, hence $x\in Fix(h^{-1}rh)\cap Fix(r^{ld})$. As before $[r,hr^{ld}h^{-1}]$ is trivial in $G$, and so is its conjugate $[h^{-1}rh,r^{ld}]$. By Lemma 3.4 applied to $Stab(x)$ and the fact that $r$ is not a proper power in $Stab(x)$, there exists $p\in\mathbb{Z}$ such that $h^{-1}rh=r^{p}$. Thus $r^{pln}=h^{-1}r^{ln}h=r^{lm}$ and $pln=lm$ because $r$ has infinite order. Dividing, we see that $lm/ln=p$ is an integer. The equation $hr^{lm/ln}h^{-1}=r^{ln/ln}$ and the fact that $r$ is very unbalanced imply that $|ln|=1$, a contradiction. ∎
Remark 3.6.
Note that without the splitting assumption, a residually finite group can indeed contain a very unbalanced element: Consider the subgroup of $GL_{2}(\mathbb{Q})$ generated by $a=\begin{pmatrix}2/3&0\\ 0&1\end{pmatrix}$ and $b=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$. As a finitely generated, linear group, it is residually finite by [Mal40], yet $a$ conjugates $b^{2}$ and $b^{3}$. This group is in fact metabelian, splitting as $\mathbb{Z}\left[\frac{1}{6}\right]$$-\mathbb{Z}$.
Lemma 3.7.
Let $A,B$ be arbitrary groups. The free product $A*B$ contains a very unbalanced element if and only if $A$ or $B$ does.
@@
Conversely, if $g\in A*B$ is very unbalanced for $h,m,n$, then $g$ is elliptic in the Bass-Serre tree of the free product, fixing a vertex $v$, as in the proof of Proposition 3.5. The element $hg^{m}h^{-1}=g^{n}$ is non-trivial (as $g$ is infinite-order) and fixes both $v$ and $hv$. Since edge stabilizers are trivial, $v=hv$, proving that $g$ is very unbalanced for $h,m,n$ inside the subgroup $Stab(v)$ which is conjugate to either $A$ or $B$. ∎
@@
When the splitting of $G$ has finite underlying graph, $F$ can be taken of finite rank, and the splitting of $G*F$ has finite underlying graph as well. In this case, the union of edge groups generate a finite-rank subgroup of $H$, which is contained in a finite-rank free factor. Thus, $H$ decomposes as a free product $K_{1}*K_{2}$ such that $K_{1}$ is finite-rank and contains all the edge groups. This induces a new splitting of $G*F$ as $G^{\prime}*K_{2}$, where $G^{\prime}$ splits as a finite multiple HNN-extension of $K_{1}$ along cyclic edge groups. Note that $F$ and $K_{2}$ are residually finite and do not contain very unbalanced elements (by Proposition 3.5).
The results of this paper state implications between residual finiteness and absence of very unbalanced elements. By Lemma 2.8 and Lemma 3.7, instead of working with some arbitrary graph of free groups with cyclic edge groups $G$ we can restrict to working with $G*F$, a multiple HNN-extension of a free group along cyclic edge groups. Moreover when the initial splitting of $G$ is finite, we can further restrict to working with $G^{\prime}$, a finite multiple HNN-extension of a finite-rank free group along cyclic edge groups.

### v2

@@ Abstract.
We characterize which groups splitting as finite graphs of free groups with cyclic edge groups are residually finite. Such a group $G$ is residually finite if and only if all its Baumslag-Solitar subgroups are residually finite. From a presentation of $G$, we construct a finite labeled graph $\Gamma$, and show that residual finiteness of $G$ is equivalent to an easily-detectable property of this graph. This characterization proves a conjecture of Wise from [Wis00].
@@ 1. Introduction
Suppose $G$ splits as a finite graph of free groups with cyclic edge groups. Then $G$ is residually finite if and only if the Baumslag-Solitar subgroups of $G$ are residually-finite.
Previously, Wise characterized the LERF (i.e.  subgroup separable) groups in this family, a much stronger property than residual finiteness [Wis00]. He proved that a group $G$ splitting as a finite graph of free groups with cyclic edge groups is LERF if and only if there exists no unbalanced element, an infinite-order $g$ such that $g^{n}$ and $g^{m}$ are conjugate for $|n|$, $|m|$ distinct. We prove (Proposition 3.6) that this is equivalent to all the Baumslag-Solitar subgroups of $G$ being LERF.
Our proof uses an analogous characterization of residual finiteness: a group splitting as a graph of free groups with cyclic edge groups is residually finite if and only if it contains no very unbalanced elements (see Definition 3.1). The existence of a very unbalanced element is easily detectable from the graph of groups splitting of $G$ (Remark 4.10). In particular, there exists an algorithm which either confirms $G$ is residually finite or outputs a very unbalanced element.
@@
In Section 3 we define and discuss very unbalanced group elements. In general, a residually finite group can contain very unbalanced elements (Remark 3.3). However, a residually finite group splitting as a graph of free groups cannot contain them, as they imply the existence of non residually finite Baumslag-Solitar subgroups (Proposition 3.6). In other words, when discussing graphs of free groups, very unbalanced elements serve as a certificate of non residual finiteness. This provides one direction of Theorem A.
@@ 2.4. Small-cancellation theory
3. Unbalanced elements and Baumslag-Solitar subgroups
@@
Let $g\in G$ be an infinite-order element in a group. Say $g$ is unbalanced for $h\in G$, $m,n\in\mathbb{Z}\setminus\{0\}$ if $|m|\neq|n|$ and $hg^{m}h^{-1}=g^{n}$ holds.
Say $g$ is very unbalanced for $h\in G$, $m,n\in\mathbb{Z}\setminus\{0\}$ if $1$, $|m|$ and $|n|$ are distinct and the following hold:
@@
  • If $n$ divides $m$, $hg^{m/n}h^{-1}\neq g$
  • If $m$ divides $n$, $hgh^{-1}\neq g^{n/m}$.
The definitions above are clearly invariant under conjugacy.
Example 3.2.
Let $p,q$ be non-zero integers. The Baumslag-Solitar group $BS(p,q)$ is defined by the presentation $\langle a,t\mid ta^{p}t^{-1}=a^{q}\rangle$. If $|p|\neq|q|$, then the element $a\in BS(p,q)$ is unbalanced. If $1$, $|p|$, and $|q|$ are distinct, then $a\in BS(p,q)$ is very unbalanced (this is a consequence of Britton’s Lemma, see Lemma 4.4).
Remark 3.3.
Wise ([Wis00]) identified the presence of unbalanced elements as an obstruction for any group to be LERF. Here we prove that very unbalanced elements are an obstruction for a group splitting as a graph of free groups to be residually finite. However, this is not true for an arbitrary group: Consider the subgroup of $GL_{2}(\mathbb{Q})$ generated by $a=\begin{pmatrix}3/2&0\\ 0&1\end{pmatrix}$ and $b=\begin{pmatrix}1&1\\ 0&1\end{pmatrix}$. As a finitely generated, linear group, it is residually finite by [Mal40], yet $b$ is very unbalanced for $a$, $2$, $3$. This group is in fact metabelian, splitting as $\mathbb{Z}\left[\frac{1}{6}\right]$$-\mathbb{Z}$.
@@
Let $G$ split as a graph of groups, with Bass-Serre tree $\mathcal{T}$, and let $g\in G$ be unbalanced for $h,m,n$. Then $g$ fixes a vertex of $\mathcal{T}$. Moreover, if the vertex groups of the splitting are free, $h$ does not fix any vertex of $\mathcal{T}$ and $g^{n}$ fixes an edge of $\mathcal{T}$.
@@
Let $d=gcd(m,n)$. Since $g^{m}$ and $g^{n}$ are conjugates in $G$, they act on $\mathcal{T}$ with the same translation length $\ell(g^{m})=\ell(g^{n})$. Then $\ell(g)=\ell(g^{m})/m=\ell(g^{n})/n=0$, i.e. $g$ fixes a vertex $v$ (see [Ser80, Proposition 25]).
The element $g^{n}=hg^{m}h^{-1}$ fixes the geodesic segment $S$ joining $v$ and $hv$. Assuming $h$ fixes some vertex of $\mathcal{T}$, $h$ fixes a vertex $w$ of $S$. Then the free group $Stab(w)$ contains $h$ and $g^{n}$, hence contains $g^{m}$ and $g^{d}\in\left\langle g^{m},g^{n}\right\rangle$. The element $g^{d}$ is unbalanced in $Stab(w)$ for $h,m/d,n/d$, hence fixes a vertex of the Cayley tree for $Stab(w)$ by the same translation length argument as before, a contradiction. Thus $h$ fixes no vertex, and $g^{n}$ fixes all the edges joining the distinct vertices $v$ and $hv$. ∎
@@
For every non-zero integer $p$, $BS(1,p)$ contains no very unbalanced elements.
@@
Using the relation $tat^{-1}=a^{p}$, every element $h\in BS(1,p)=\left\langle a,t\right\rangle$ can be represented as a product of the form $t^{-i}a^{j}t^{k}$ with $i,k\geq 0$ and $j\in\mathbb{Z}$.
Assume $g\in BS(1,p)$ is very unbalanced for $h,m,n$. By Lemma 3.4, $g$ fixes a vertex of the Bass-Serre tree of the HNN extension defining $BS(1,p)$. Up to conjugating, assume $g$ is of the form $a^{l}$, $l\neq 0$. Let $i,k\geq 0$, $j\in\mathbb{Z}$ so that $h=t^{-i}a^{j}t^{k}$. Then the following equivalent equations hold:
$g^{n}$ $=hg^{m}h^{-1}$
$a^{nl}$ $=t^{-i}a^{j}t^{k}a^{ml}t^{-k}a^{-j}t^{i}$
$t^{i}a^{nl}t^{-i}$ $=a^{j}t^{k}a^{ml}t^{-k}a^{-j}$
$a^{p^{i}nl}$ $=a^{j}a^{p^{k}ml}a^{-j}=a^{p^{k}ml}$
Therefore $p^{i}nl=p^{k}ml$, since $a$ has infinite order. If $i<k$, then $n=p^{k-i}m$ is a multiple of $m$ and $hgh^{-1}=t^{-i}a^{p^{k}l}t^{i}=a^{p^{k-i}l}=g^{n/m}$ holds. Symmetrically if $i>k$, then $m$ is a multiple of $n$ and $hg^{m/n}h^{-1}=g$ holds. If $i=k$, then $m=n$. This contradicts the fact that $g$ is very unbalanced. ∎
For the next result, recall (see [Mol18]) that $BS(p,q)$ is LERF if and only if $|p|=|q|$; $BS(p,q)$ is residually finite if and only if $1$, $|p|$, $|q|$ are not distinct; and both being LERF and residually finite are preserved under passing to a subgroup.
Proposition 3.6.
Let $G$ split as a graph of free groups.
  1. (1)
    $G$ contains an unbalanced element if and only if there exists $p,q\in\mathbb{Z}\setminus\{0\}$ with $|p|\neq|q|$ such that $BS(p,q)$ embeds in $G$. In that case, $G$ is not LERF.
  2. (2)
    $G$ contains a very unbalanced element if and only if there exists $p,q\in\mathbb{Z}\setminus\{0\}$ with $1$, $|p|$ and $|q|$ distinct such that $BS(p,q)$ embeds in $G$. In that case, $G$ is not residually finite.
Proof.
The "if" directions of both statements are clear since an unbalanced (resp. very unbalanced) element of a subgroup of $G$ is unbalanced (resp. very unbalanced) in $G$ as well.
For the "only if" direction, let $g\in G$ be unbalanced for $h,m,n$ and let $d=gcd(m,n)$. Let $\mathcal{T}$ be the Bass-Serre tree of the splitting of $G$ and $\mathcal{T}^{\prime}$ the subtree given by the union $\bigcup_{k\geq 1}Fix(g^{k})$. The union is connected since all the terms contain $Fix(g)$. For every vertex $v$ of $\mathcal{T}^{\prime}$, there exists $k\geq 1$ such that $v\in Fix(g^{k})\subseteq Fix(g^{mk})$. Thus $hv\in Fix(hg^{mk}h^{-1})=Fix(g^{nk})\subseteq\mathcal{T}^{\prime}$. Therefore, $\mathcal{T}^{\prime}$ is $\left\langle g,h\right\rangle$-stable.
Moreover, for every $h^{\prime}\in\left\langle g,h\right\rangle\cap Stab(v)$, $h^{\prime}$ conjugates $g^{km^{K}n^{K}}$ to $g^{kL}$, for some integers $K$ and $L$, when $K$ is sufficiently large. Note that $h^{\prime}$, $g^{km^{K}n^{K}}$ and $g^{kL}$ are in $Stab(v)$. Since $Stab(v)$ is free, $g^{km^{K}n^{K}}$ and $g^{kL}$ have the same translation length in the $Stab(v)$ Cayley tree, hence $|km^{K}n^{K}|=|kL|$. Therefore, $h^{\prime 2}$ and $g^{kL}$ commute in $Stab(v)$, meaning that $h^{\prime}$ is contained in the cyclic subgroup generated by a maximal root $r$ of $g^{kL}$ in $Stab(v)$. Since $\left\langle g^{kL}\right\rangle\subseteq\left\langle g,h\right\rangle\cap Stab(v)\subseteq\left\langle r\right\rangle$, the group $\left\langle g,h\right\rangle\cap Stab(v)$ is infinite cyclic. This holds for any vertex $v$ of $\mathcal{T}^{\prime}$. Moreover, for an arbitrary edge $e$ of $\mathcal{T}^{\prime}$, $\left\langle g,h\right\rangle\cap Stab(e)$ contains a non-trivial power of $g$ and is contained in a vertex stabilizer, thus $\left\langle g,h\right\rangle\cap Stab(e)$ is infinite cyclic as well.
The action of $\left\langle g,h\right\rangle$ on $\mathcal{T}^{\prime}$ has infinite cyclic vertex and edges stabilizers, and therefore provides a splitting of $\left\langle g,h\right\rangle$ as a Generalized Baumslag-Solitar group. Since $hg^{m}h^{-1}=g^{n}$ and $g$ has infinite order, $\left\langle g,h\right\rangle$ is a non-abelian quotient of $BS(m,n)$.
By the argument above applied to $g^{d}$, which is unbalanced for $h,m/d,n/d$, $\left\langle g^{d},h\right\rangle$ splits as a Generalized Baumslag-Solitar group and is a non-abelian quotient of $BS(m/d,n/d)$. Moreover, $|m/d|$ and $|n/d|$ are coprime. By a theorem of Levitt ([Lev15, Lemma 7.4]), $BS(m/d,n/d)$ embeds in $\left\langle g^{d},h\right\rangle$, proving Assertion (1).
Now assume further that $g$ is very unbalanced for $h$, $m$, $n$. If $m$ is not a multiple or a divisor of $n$, then $1$, $|m/d|$ and $|n/d|$ are distinct, and the previous argument suffices to prove Assertion (2) as well. Otherwise, say $m$ divides $n$, the other case being symmetric. By Lemma 3.5, the Generalized Baumslag-Solitar group $\left\langle g,h\right\rangle$ is not isomorphic to $BS(1,p)$ for any $p$, i.e. not solvable. Yet $BS(1,n/m)$ embeds in $\left\langle g^{m},h\right\rangle\leq\left\langle g,h\right\rangle$ by the argument above. By another theorem of Levitt ([Lev15, Lemma 7.6]), there exists a prime $q$ such that $BS(q,qn/m)$ embeds in $\left\langle g,h\right\rangle$, proving Assertion (2) in that case. ∎
Remark 3.7.
A variant of the previous argument justifies that when $g\in G$ is conjugate into an edge group, the commensurator of $\left\langle g\right\rangle$ in $G$ splits as a Generalized Baumslag-Solitar group. By [Wis00], the absence of unbalanced elements in $G$ is equivalent to all such commensurators being LERF (i.e. unimodular in Levitt’s terminology). When the edge groups are cyclic, it is also equivalent to $G$ itself being LERF.
Likewise, our main result implies that the absence of very unbalanced elements in $G$ is equivalent to all such commensurators being residually finite (i.e. unimodular or isomorphic to some $BS(1,p)$). When the edge groups are cyclic, it is also equivalent to $G$ itself being residually finite.
@@
Let $A,B$ be arbitrary groups. The free product $A*B$ contains a very unbalanced element if and only if $A$ or $B$ does. Moreover, for $|p|\neq|q|$, $BS(p,q)$ embeds in $A*B$ if and only if it embeds in $A$ or in $B$
@@
Conversely, if $g\in A*B$ is very unbalanced for $h,m,n$, then $g$ fixes a vertex in the Bass-Serre tree of the free product, fixing a vertex $v$, by Lemma 3.4. The element $hg^{m}h^{-1}=g^{n}$ is non-trivial (as $g$ is infinite-order) and fixes both $v$ and $hv$. Since edge stabilizers are trivial, $v=hv$, proving that $g$ is very unbalanced for $h,m,n$ inside the subgroup $Stab(v)$ which is conjugate to either $A$ or $B$.
Moreover, if $BS(p,q)$ embeds in $A$ or $B$, it clearly embeds in $A*B$ through the monomorphisms $A\to A*B$, $B\to A*B$. Conversely, if $BS(p,q)=\left\langle a,t\right\rangle$ embeds in $A*B$, $a$ fixes a vertex $v$ in the Bass-Serre tree of the free product by Lemma 3.4. Since edge stabilizers are trivial and $a$ has infinite order, $Fix(a^{p})=Fix(a^{q})=\{v\}$. However, $\{v\}=Fix(a^{q})=Fix(ta^{p}t^{-1})=tFix(a^{p})=\{tv\}$, hence $\left\langle a,t\right\rangle\leq Stab(v)$. Up to conjugating, $BS(p,q)=\left\langle a,t\right\rangle$ embeds in $A$ or $B$. ∎
@@
When the splitting of $G$ has finite underlying graph, $F$ can be taken of finite rank, and the splitting of $G*F$ has finite underlying graph as well. In this case, the union of edge groups generate a finite-rank subgroup of $H$, which is contained in a finite-rank free factor. Thus, $H$ decomposes as a free product $K_{1}*K_{2}$ such that $K_{1}$ is finite-rank and contains all the edge groups. This induces a new splitting of $G*F$ as $G^{\prime}*K_{2}$, where $G^{\prime}$ splits as a finite multiple HNN-extension of $K_{1}$ along cyclic edge groups. Note that $F$ and $K_{2}$ are residually finite and do not contain very unbalanced elements nor $BS(p,q)$ subgroups.
The results of this paper state implications between residual finiteness, absence of very unbalanced elements and absence of certain Baumslag-Solitar subgroups. By Lemma 2.8 and Lemma 3.8, instead of working with some arbitrary graph of free groups with cyclic edge groups $G$ we can restrict to working with $G*F$, a multiple HNN-extension of a free group along cyclic edge groups. Moreover when the initial splitting of $G$ is finite, we can further restrict to working with $G^{\prime}$, a finite multiple HNN-extension of a finite-rank free group along cyclic edge groups.